# Breaking the Dimensional Barrier: A Pontryagin-Guided Direct Policy Optimization for Continuous-Time Multi-Asset Portfolio

### v3

@@ 2.3.1 A Multi-Asset Model with One OU Factor
We first consider a model with $n$ risky assets whose risk premia are driven by a single Ornstein–Uhlenbeck (OU) factor. Specifically, each asset $i$ has a constant volatility $\sigma_{i}>0$ and a time-varying risk premium proportional to an OU state variable $Y_{t}$, such that the expected return is $r+\alpha_{i}\sigma_{i}Y_{t}$, where $r$ is the constant risk-free rate. The state variable $Y_{t}$ follows the OU process:
@@
where $\kappa_{Y}>0$ is the mean-reversion speed, $\theta_{Y}\in\mathbb{R}$ is the long-term mean, and $\sigma_{Y}>0$ is the factor volatility. The Brownian motion $W_{t}^{Y}$ may be correlated with the asset return processes.
The price dynamics for the $i$-th risky asset, $S_{i,t}$, are given by:
$\frac{dS_{i,t}}{S_{i,t}}=(r+\alpha_{i}\sigma_{i}Y_{t})\,dt+\sigma_{i}\,dW_{i,t}^{S}.$
Let $\bm{\alpha}=(\alpha_{1},\dots,\alpha_{n})^{\top}$ be the vector of loading coefficients and $\bm{\sigma}=\operatorname{diag}(\sigma_{1},\dots,\sigma_{n})$ be the diagonal matrix of constant asset volatilities, where each $\sigma_{i}>0$. The Brownian motions driving the assets, $\mathbf{W}_{t}^{S}=(W_{1,t}^{S},\dots,W_{n,t}^{S})^{\top}$, have an $n\times n$ covariance matrix $\Psi dt$, i.e., $\mathrm{Cov}(d\mathbf{W}_{t}^{S})=\Psi\,dt$. Furthermore, each $dW_{i,t}^{S}$ may have a correlation $\rho_{i}$ with $dW_{t}^{Y}$; let $\bm{\rho}=(\rho_{1},\dots,\rho_{n})^{\top}$.
Let $X_{t}$ denote total wealth. The investor chooses a consumption rate $C_{t}\geq 0$ and portfolio weights $\bm{\pi}_{t}\in\mathbb{R}^{n}$ for the risky assets. The wealth dynamics follow:
@@
The instantaneous covariance matrix of the $n$-asset returns is $\Sigma=\bm{\sigma}\,\Psi\,\bm{\sigma}$.
For an investor with CRRA utility $U(x)$ and discount rate $\delta>0$, the optimal consumption $C_{t}^{*}$ is determined by $C_{t}^{*}=(U^{\prime})^{-1}(e^{\delta t}\lambda_{t}^{*})$, where $\lambda_{t}^{*}$ is the costate associated with wealth $X_{t}$. The PMP first-order condition yields the optimal portfolio $\bm{\pi}_{t}^{*}\in\mathbb{R}^{n}$ as:
$\bm{\pi}_{t}^{*}=-\frac{1}{X_{t}^{*}(\partial_{x}\lambda_{t}^{*})}\Sigma^{-1}\left\{\lambda_{t}^{*}Y_{t}(\bm{\sigma}\bm{\alpha})+(\partial_{Y}\lambda_{t}^{*})\sigma_{Y}(\bm{\sigma}\bm{\rho})\right\}.$
This expression involves the costate $\lambda_{t}^{*}$ and its derivatives $\partial_{x}\lambda_{t}^{*}$ and $\partial_{Y}\lambda_{t}^{*}$. The first term within the braces represents the myopic demand driven by the risk premium $Y_{t}(\bm{\sigma}\bm{\alpha})$, and the second term captures the intertemporal hedging demand arising from the state variable $Y_{t}$. If $n=1$, this simplifies to $\pi_{t}^{*}=-\frac{1}{\sigma^{2}X_{t}^{*}\partial_{x}\lambda_{t}^{*}}\{\lambda_{t}^{*}\alpha\sigma Y_{t}+\sigma(\partial_{Y}\lambda_{t}^{*})\sigma_{Y}\rho\}$. For the specific case without intermediate consumption ($C_{t}=0$), an analytic solution involving Riccati ODEs can be derived using the HJB framework (Kim and Omberg, , 1996), as detailed in Appendix B.1.
@@ 2.3.2 A Multi-Asset, Multi-Factor Model
We then extend this setup to a more general multi-factor environment, considering $n$ risky assets whose risk premia are driven by a $k$-dimensional OU factor vector $\mathbf{Y}_{t}\in\mathbb{R}^{k}$. The risk premium for asset $i=1,\dots,n$ is $\sigma_{i}(\bm{\alpha}_{i}^{\top}\mathbf{Y}_{t})$, where $\bm{\alpha}_{i}\in\mathbb{R}^{k}$ contains the factor loadings for asset $i$. The volatility $\sigma_{i}$ for asset $i$ is constant. This allows for richer interactions between multiple risk sources and asset returns.
The factor vector $\mathbf{Y}_{t}\in\mathbb{R}^{k}$ follows a multivariate OU process:
$d\mathbf{Y}_{t}\;=\;\kappa_{Y}\bigl{(}\bm{\theta}_{Y}-\mathbf{Y}_{t}\bigr{)}\,dt\;+\;\bm{\sigma}_{Y}\,d\mathbf{W}_{t}^{Y},$
@@
where $\kappa_{Y}\in\mathbb{R}^{k\times k}$ is a mean-reversion matrix, $\bm{\theta}_{Y}\in\mathbb{R}^{k}$ is the long-run mean vector, $\bm{\sigma}_{Y}$ is the $k\times k$ diagonal matrix of factor-specific diffusion coefficients (as defined in Section 2.1), and $d\mathbf{W}_{t}^{Y}$ is a $k$-dimensional vector of Brownian motions whose components may be correlated, with $\mathrm{Cov}(d\mathbf{W}_{t}^{Y})=\Phi dt$.
The $n$ risky assets, with prices $S_{t}^{i}$, follow:
@@
$\frac{dS_{t}^{i}}{S_{t}^{i}}\;=\;\Bigl{(}r+\sigma_{i}\,\bm{\alpha}_{i}^{\top}\mathbf{Y}_{t}\Bigr{)}\,dt\;+\;\sigma_{i}\,dW_{t}^{X,i},\quad i=1,\dots,n.$
@@
Here, $\mathbf{W}_{t}^{X}\in\mathbb{R}^{n}$ is a Brownian motion with covariance matrix $\Psi dt$. The asset shocks $\mathbf{W}_{t}^{X}$ and factor shocks $\mathbf{W}_{t}^{Y}$ can be correlated, with $\mathrm{Cov}(dW_{t}^{X,i},dW_{t}^{Y,j})=\rho_{ij}dt$. Let $\bm{\rho}\in\mathbb{R}^{n\times k}$ be the matrix of cross-correlations.
@@
With portfolio weights $\bm{\pi}_{t}=(\pi_{t}^{1},\dots,\pi_{t}^{n})^{\top}$ and consumption $C_{t}$, wealth $X_{t}$ evolves as:
$dX_{t}=\left[X_{t}\left(r+\sum_{i=1}^{n}\pi_{t}^{i}\sigma_{i}(\bm{\alpha}_{i}^{\top}\mathbf{Y}_{t})\right)-C_{t}\right]dt+X_{t}\sum_{i=1}^{n}\pi_{t}^{i}\sigma_{i}\,dW_{t}^{X,i}.$
@@
For CRRA utility with discount $\delta$, the optimal consumption $C_{t}^{*}$ depends on the wealth costate $\lambda_{t}^{*}$. The optimal portfolio $\bm{\pi}_{t}^{*}\in\mathbb{R}^{n}$ is given by the PMP first-order condition (Eq. (11)), which in this context becomes:
$\bm{\pi}_{t}^{*}=-\;\frac{1}{\,X_{t}^{*}\,\bigl{(}\partial_{x}\lambda_{t}^{*}\bigr{)}\!}\;\Sigma^{-1}\;\Bigl{\{}\,\lambda_{t}^{*}\;\bm{\sigma}\,\mathbf{A}\,\mathbf{Y}_{t}~{}+~{}\bm{\sigma}\,\bm{\rho}\,\bm{\sigma}_{Y}\partial_{\mathbf{Y}}\lambda_{t}^{*}\Bigr{\}},$
@@
where $\bm{\sigma}=\operatorname{diag}(\sigma_{1},\dots,\sigma_{n})$ is the $n\times n$ diagonal matrix of asset volatilities, $\Sigma=\bm{\sigma}\Psi\bm{\sigma}$ is the $n\times n$ asset return covariance, $\mathbf{A}=(\bm{\alpha}_{1},\dots,\bm{\alpha}_{n})^{\top}$ is the $n\times k$ matrix of factor loadings (where each $\bm{\alpha}_{i}$ is a $k\times 1$ vector of loadings for asset $i$), $\bm{\sigma}_{Y}$ is the $k\times k$ diagonal matrix of factor diffusion coefficients, and $\partial_{\mathbf{Y}}\lambda_{t}^{*}$ is the $k\times 1$ gradient vector of the costate with respect to the factors $\mathbf{Y}_{t}$. This explicitly separates the demand into a myopic part driven by the current expected excess returns ($\bm{\sigma}\,\mathbf{A}\,\mathbf{Y}_{t}$) and a hedging part driven by the sensitivity to factor changes ($\partial_{\mathbf{Y}}\lambda_{t}^{*}$) and their correlations ($\bm{\rho}$, $\Psi$, $\Phi$).
Similar to the single-factor case, an analytic solution for the terminal wealth problem ($C_{t}=0$) with CRRA utility can often be found using an exponential-quadratic value function ansatz, leading to matrix Riccati ODEs. Details are presented in Appendix B.2.
@@ 3 A Gradient-Based Algorithm for Multi-Asset Policy Optimization
While Pontryagin’s Maximum Principle (PMP) provides necessary optimality conditions via forward-backward stochastic differential equations (FBSDEs, Section 2.2), solving these analytically is often infeasible in high dimensions. We therefore adopt a gradient-based numerical approach using deep learning. The core idea is to parameterize the investor’s policy $(\bm{\pi}_{t},C_{t})$ with neural networks $(\bm{\pi}_{\theta}(t,x,\mathbf{y}),C_{\phi}(t,x,\mathbf{y}))$, where $(\theta,\phi)$ are the parameters to be optimized.
The objective is to find parameters $(\theta,\phi)$ that maximize the expected utility achieved by the policy $(\bm{\pi}_{\theta},C_{\phi})$. To ensure robustness across different starting scenarios, we aim to maximize the extended value function, $\widetilde{J}(\theta,\phi)$. This is defined as the expected value of the conditional expected utility $J(t_{0},x_{0},\mathbf{y}_{0};\theta,\phi)$ achieved under policy $(\theta,\phi)$, where the initial state $(t_{0},x_{0},\mathbf{y}_{0})$ is drawn from a distribution $\eta$ over a relevant domain $\mathcal{D}$:
@@
$\widetilde{J}(\theta,\phi)=\mathbb{E}_{(t_{0},x_{0},\mathbf{y}_{0})\,\sim\,\eta}\Big{[}J(t_{0},x_{0},\mathbf{y}_{0};\theta,\phi)\Big{]}$ (13)
Here, $\mathbb{E}^{\theta,\phi}[\cdot|\cdot]$ denotes expectation under the dynamics generated by policy $(\theta,\phi)$.
This paper develops and implements the Pontryagin-Guided Direct Policy Optimization (PG-DPO) framework to maximize $\widetilde{J}(\theta,\phi)$. PG-DPO leverages automatic differentiation, specifically backpropagation-through-time (BPTT), applied to Monte Carlo simulations drawn from the initial state distribution $\eta$. A key feature is that BPTT applied to these simulations yields not only unbiased estimates of the policy gradient $\nabla\widetilde{J}$ needed for optimization, but also inherently calculates estimates of the Pontryagin costate variables ($\lambda_{k}^{(i)},\bm{\eta}_{k}^{(i)}$) along each path, representing sensitivities dictated by PMP.
This dual output enables distinct algorithmic approaches: the baseline PG-DPO method is detailed in Section 3.1, and the advanced “Two-Stage PG-DPO” variant, which leverages these costate estimates more directly, is explored in Section 3.2.
@@ 3.1 Baseline PG-DPO for the Extended Objective
    where $\Delta t^{(i)}=(T-t_{0}^{(i)})/N$, $\bm{\sigma}_{Y}(t_{k}^{(i)},\mathbf{Y}_{k}^{(i)})$ is the vector of diffusion coefficients applied element-wise to the components of $\Delta\mathbf{W}_{k}^{(i),Y}$, and the correctly correlated Brownian increments $(\Delta\mathbf{W}_{k}^{(i),X},\Delta\mathbf{W}_{k}^{(i),Y})$ are derived from the Cholesky factor $L$ of the overall covariance matrix $\bm{\Omega}$.
@@
It is important to understand the role of the intermediate values computed during BPTT in Step 4. As BPTT calculates the policy gradient $\nabla_{(\theta,\phi)}J^{(i)}(\dots)$ via the chain rule, it inherently computes the partial derivatives of the realized reward $J^{(i)}(\dots)$ with respect to the intermediate states along path $i$:
$\lambda_{k}^{(i)}=\frac{\partial J^{(i)}(t_{0}^{(i)},\dots)}{\partial X_{k}^{(i)}},\quad\bm{\eta}_{k}^{(i)}=\frac{\partial J^{(i)}(t_{0}^{(i)},\dots)}{\partial\mathbf{Y}_{k}^{(i)}}.$
These sensitivities, computed mechanically by AD, are directly linked to the discrete adjoint Backward Stochastic Differential Equation (BSDE) structure required by Pontryagin’s Maximum Principle (PMP). To illustrate this crucial connection for the wealth component (denoting $\lambda_{k}^{(i)}$ for $\lambda_{k}^{X,(i)}$ for simplicity), consider the discrete Euler-Maruyama step for wealth under policy $(\bm{\pi}_{k}^{(i)},C_{k}^{(i)})$:
@@
$X_{k+1}^{(i)}=X_{k}^{(i)}+\Bigl{[}X_{k}^{(i)}\Bigl{(}r(t_{k},\mathbf{Y}_{k}^{(i)})+\bm{\pi}_{k}^{(i)\top}\bigl{(}\bm{\mu}(t_{k},\mathbf{Y}_{k}^{(i)})-r(t_{k},\mathbf{Y}_{k}^{(i)})\mathbf{1}\bigr{)}\Bigr{)}-C_{k}^{(i)}\Bigr{]}\Delta t+X_{k}^{(i)}\bm{\pi}_{k}^{(i)\top}\bm{\sigma}(t_{k},\mathbf{Y}_{k}^{(i)})\,\Delta\mathbf{W}_{k}^{(i)}.$
Applying the chain rule one step backward ($\lambda_{k}^{(i)}=\lambda_{k+1}^{(i)}\frac{\partial X_{k+1}^{(i)}}{\partial X_{k}^{(i)}}$) and using a first-order expansion yields the following backward recurrence relation along the path $(i)$ (this illustrative recurrence primarily shows influences from the wealth dynamics; a full derivation of the drift for $\lambda_{k}^{(i)}$ from the total reward $J^{(i)}$ would also account for terms related to consumption $U(C_{k}^{(i)})$ and the discount factor in its definition within the Hamiltonian context):

### v4

@@ 2.3.1 A Multi-Asset Model with One OU Factor
We first consider a model with $n$ risky assets whose excess returns are driven by a single Ornstein–Uhlenbeck (OU) factor. Specifically, each asset $i$ has constant volatility $\sigma_{i}>0$ and a time-varying risk premium proportional to an OU state variable $Y_{t}$, so that the expected return is $r+\alpha_{i}\sigma_{i}Y_{t}$, where $r$ is the constant risk-free rate. The state variable $Y_{t}$ evolves as:
@@
where $\kappa_{Y}>0$ is the mean-reversion speed, $\theta_{Y}\in\mathbb{R}$ is the long-run mean, and $\sigma_{Y}>0$ is the volatility of the OU process. The Brownian motion $W_{t}^{Y}$ may be correlated with the return processes of the risky assets.
The price process $S_{i,t}$ of the $i$-th risky asset follows:
$\frac{dS_{i,t}}{S_{i,t}}=\bigl{(}r+\alpha_{i}\sigma_{i}Y_{t}\bigr{)}\,dt+\sigma_{i}\,dW_{i,t}^{S}.$
Let $\bm{\alpha}=(\alpha_{1},\dots,\alpha_{n})^{\top}$ be the vector of risk premia loadings, and $\bm{\sigma}=\operatorname{diag}(\sigma_{1},\dots,\sigma_{n})$ the diagonal matrix of asset volatilities. The Brownian motion driving the risky assets, $\mathbf{W}_{t}^{S}=(W_{1,t}^{S},\dots,W_{n,t}^{S})^{\top}$, has instantaneous covariance matrix $\mathrm{Cov}(d\mathbf{W}_{t}^{S})=\Psi\,dt$, where $\Psi\in\mathbb{R}^{n\times n}$. Let $\bm{\rho}=(\rho_{1},\dots,\rho_{n})^{\top}$ denote the vector of correlations between $W_{t}^{Y}$ and $W_{i,t}^{S}$.
Let $X_{t}$ denote the investor’s wealth. The investor chooses a consumption rate $C_{t}\geq 0$ and portfolio weights $\bm{\pi}_{t}\in\mathbb{R}^{n}$ in the risky assets. The wealth dynamics is given by:
@@
The instantaneous covariance matrix of risky asset returns is $\Sigma=\bm{\sigma}\Psi\bm{\sigma}$.
For an investor with CRRA utility $U(x)$ and discount rate $\delta>0$, the optimal consumption is given by $C_{t}^{*}=(U^{\prime})^{-1}(e^{\delta t}\lambda_{t}^{*})$, where $\lambda_{t}^{*}$ is the costate variable associated with wealth $X_{t}$. The first-order condition from Pontryagin’s Maximum Principle yields the optimal portfolio:
$\bm{\pi}_{t}^{*}=-\frac{1}{X_{t}^{*}\,\partial_{x}\lambda_{t}^{*}}\,\Sigma^{-1}\left\{\lambda_{t}^{*}Y_{t}(\bm{\sigma}\bm{\alpha})+(\partial_{Y}\lambda_{t}^{*})\,\sigma_{Y}\,(\bm{\sigma}\bm{\rho})\right\}.$
This expression depends on the costate $\lambda_{t}^{*}$ and its derivatives $\partial_{x}\lambda_{t}^{*}$ and $\partial_{Y}\lambda_{t}^{*}$. The first term inside the braces represents the myopic demand, driven by the instantaneous risk premia $Y_{t}(\bm{\sigma}\bm{\alpha})$. The second term captures the intertemporal hedging demand, which arises from uncertainty in future investment opportunities and is scaled by the correlation vector $\bm{\rho}$ and the sensitivity $\partial_{Y}\lambda_{t}^{*}$.
In the special case of a single risky asset ($n=1$), the portfolio simplifies to:
$\pi_{t}^{*}=-\frac{1}{\sigma^{2}X_{t}^{*}\,\partial_{x}\lambda_{t}^{*}}\left\{\lambda_{t}^{*}\alpha\sigma Y_{t}+\sigma\rho\sigma_{Y}\,\partial_{Y}\lambda_{t}^{*}\right\}.$
In the case without intermediate consumption ($C_{t}\equiv 0$), the value function and optimal controls can be derived analytically via the Hamilton–Jacobi–Bellman (HJB) approach, leading to a system of Riccati ODEs (Kim and Omberg, , 1996). Details are provided in Appendix B.1.
@@ 2.3.2 A Multi-Asset, Multi-Factor Model
We next extend the previous setup to a more general multi-factor environment, considering $n$ risky assets whose excess returns are driven by a $k$-dimensional Ornstein–Uhlenbeck (OU) factor vector $\mathbf{Y}_{t}\in\mathbb{R}^{k}$. The risk premium for asset $i=1,\dots,n$ is given by $\sigma_{i}(\bm{\alpha}_{i}^{\top}\mathbf{Y}_{t})$, where $\bm{\alpha}_{i}\in\mathbb{R}^{k}$ is the vector of factor loadings for asset $i$, and $\sigma_{i}>0$ is the asset’s constant volatility. This formulation captures richer interactions between multiple sources of risk and asset returns.
The factor process $\mathbf{Y}_{t}$ evolves as a multivariate OU process:
$d\mathbf{Y}_{t}=\kappa_{Y}(\bm{\theta}_{Y}-\mathbf{Y}_{t})\,dt+\bm{\sigma}_{Y}\,d\mathbf{W}_{t}^{Y},$
@@
where $\kappa_{Y}\in\mathbb{R}^{k\times k}$ is the mean-reversion matrix, $\bm{\theta}_{Y}\in\mathbb{R}^{k}$ is the long-run mean, and $\bm{\sigma}_{Y}\in\mathbb{R}^{k\times k}$ is a diagonal matrix of factor-specific volatilities (as introduced in Section 2.1). The Brownian motion $\mathbf{W}_{t}^{Y}\in\mathbb{R}^{k}$ driving the factors has instantaneous covariance $\mathrm{Cov}(d\mathbf{W}_{t}^{Y})=\Phi\,dt$.
The price $S_{t}^{i}$ of each asset follows:
@@
$\frac{dS_{t}^{i}}{S_{t}^{i}}=\bigl{(}r+\sigma_{i}\,\bm{\alpha}_{i}^{\top}\mathbf{Y}_{t}\bigr{)}\,dt+\sigma_{i}\,dW_{t}^{X,i},\quad i=1,\dots,n,$
@@
where $\mathbf{W}_{t}^{X}\in\mathbb{R}^{n}$ is a Brownian motion with instantaneous covariance $\mathrm{Cov}(d\mathbf{W}_{t}^{X})=\Psi\,dt$. The asset and factor Brownian motions may be correlated: $\mathrm{Cov}(dW_{t}^{X,i},dW_{t}^{Y,j})=\rho_{ij}\,dt$, with cross-correlation matrix $\bm{\rho}\in\mathbb{R}^{n\times k}$.
@@
Given portfolio weights $\bm{\pi}_{t}=(\pi_{t}^{1},\dots,\pi_{t}^{n})^{\top}$ and consumption rate $C_{t}\geq 0$, the investor’s wealth evolves as:
$dX_{t}=\left[X_{t}\left(r+\sum_{i=1}^{n}\pi_{t}^{i}\sigma_{i}\bm{\alpha}_{i}^{\top}\mathbf{Y}_{t}\right)-C_{t}\right]dt+X_{t}\sum_{i=1}^{n}\pi_{t}^{i}\sigma_{i}\,dW_{t}^{X,i}.$
@@
For CRRA utility with subjective discount rate $\delta>0$, the optimal consumption policy satisfies $C_{t}^{*}=(U^{\prime})^{-1}(e^{\delta t}\lambda_{t}^{*})$, where $\lambda_{t}^{*}$ is the costate associated with wealth. The first-order condition from Pontryagin’s Maximum Principle yields the optimal portfolio $\bm{\pi}_{t}^{*}\in\mathbb{R}^{n}$ as:
$\bm{\pi}_{t}^{*}=-\;\frac{1}{X_{t}^{*}\,\partial_{x}\lambda_{t}^{*}}\,\Sigma^{-1}\left\{\lambda_{t}^{*}\,\bm{\sigma}\mathbf{A}\mathbf{Y}_{t}+\bm{\sigma}\bm{\rho}\bm{\sigma}_{Y}\,\partial_{\mathbf{Y}}\lambda_{t}^{*}\right\},$
@@
where: $\bm{\sigma}=\operatorname{diag}(\sigma_{1},\dots,\sigma_{n})\in\mathbb{R}^{n\times n}$ is the volatility matrix, $\Sigma=\bm{\sigma}\Psi\bm{\sigma}\in\mathbb{R}^{n\times n}$ is the return covariance matrix, $\mathbf{A}=(\bm{\alpha}_{1},\dots,\bm{\alpha}_{n})^{\top}\in\mathbb{R}^{n\times k}$ is the matrix of factor loadings, $\bm{\sigma}_{Y}\in\mathbb{R}^{k\times k}$ is the factor volatility matrix, $\partial_{\mathbf{Y}}\lambda_{t}^{*}\in\mathbb{R}^{k}$ is the gradient of the costate with respect to $\mathbf{Y}_{t}$.
This expression separates the optimal portfolio into two components: a myopic demand, driven by the instantaneous excess returns $\bm{\sigma}\mathbf{A}\mathbf{Y}_{t}$, and an intertemporal hedging demand, which offsets the effects of changes in investment opportunities. The latter depends on the sensitivity of marginal utility to changes in $\mathbf{Y}_{t}$ ($\partial_{\mathbf{Y}}\lambda_{t}^{*}$), and the correlation structure between asset and factor shocks (encoded in $\bm{\rho}$, $\Psi$, and $\Phi$).
As in the single-factor case, when there is no intermediate consumption ($C_{t}\equiv 0$), one can derive analytic solutions using an exponential-quadratic value function ansatz, leading to a system of matrix Riccati ODEs. Detailed derivations are provided in Appendix B.2.
@@ 3 A Gradient-Based Algorithm for Multi-Asset Policy Optimization
While Pontryagin’s Maximum Principle (PMP) provides necessary optimality conditions via a system of forward–backward stochastic differential equations (FBSDEs; see Section 2.2), solving these analytically becomes infeasible in high dimensions. We therefore adopt a gradient-based numerical approach using deep learning. The core idea is to parameterize the investor’s policy $(\bm{\pi}_{t},C_{t})$ with neural networks $(\bm{\pi}_{\theta}(t,x,\mathbf{y}),C_{\phi}(t,x,\mathbf{y}))$, where $(\theta,\phi)$ denote the network parameters to be optimized.
Our objective is to find parameters $(\theta,\phi)$ that maximize the expected utility delivered by the policy $(\bm{\pi}_{\theta},C_{\phi})$. To ensure robustness across a range of initial conditions, we define and maximize the extended value function, $\widetilde{J}(\theta,\phi)$, as the expected value of the conditional utility $J(t_{0},x_{0},\mathbf{y}_{0};\theta,\phi)$, evaluated over initial states sampled from a distribution $\eta$ on a domain $\mathcal{D}$:
@@
$\widetilde{J}(\theta,\phi)=\mathbb{E}_{(t_{0},x_{0},\mathbf{y}_{0})\,\sim\,\eta}\left[J(t_{0},x_{0},\mathbf{y}_{0};\theta,\phi)\right]$ (13)
Here, $\mathbb{E}^{\theta,\phi}[\cdot\,|\,\cdot]$ denotes expectation under the dynamics induced by policy $(\theta,\phi)$.
We develop and implement the Pontryagin-Guided Direct Policy Optimization (PG-DPO) framework to maximize $\widetilde{J}(\theta,\phi)$. PG-DPO employs automatic differentiation—specifically, backpropagation through time (BPTT)—applied to Monte Carlo simulations of system trajectories initialized from $\eta$. A key feature of this approach is that BPTT not only provides unbiased gradient estimates $\nabla\widetilde{J}$ for policy optimization, but also computes pathwise estimates of the Pontryagin costate variables $(\lambda_{k}^{(i)},\bm{\eta}_{k}^{(i)})$, capturing marginal value sensitivities as prescribed by PMP.
This dual output enables two algorithmic variants: the baseline PG-DPO algorithm described in Section 3.1, and an advanced Two-Stage PG-DPO algorithm, which explicitly exploits costate information, presented in Section 3.2.
@@ 3.1 Baseline PG-DPO for the Extended Objective
    where $\Delta t^{(i)}=(T-t_{0}^{(i)})/N$, $\bm{\sigma}_{Y}(t_{k}^{(i)},\mathbf{Y}_{k}^{(i)})$ is the vector of diffusion coefficients applied element-wise to the components of $\Delta\mathbf{W}_{k}^{(i),Y}$, and the correlated Brownian increments $(\Delta\mathbf{W}_{k}^{(i),X},\Delta\mathbf{W}_{k}^{(i),Y})$ are derived from the Cholesky factor $L$ of the covariance matrix $\bm{\Omega}$.
@@
It is important to understand the role of the intermediate values computed during BPTT. By applying the chain rule backward from the final reward $J^{(i)}$, BPTT inherently calculates the sensitivity of the reward to the intermediate state variables. These sensitivities are the pathwise numerical counterparts to the Pontryagin costates:
$\lambda_{k}^{(i)}=\frac{\partial J^{(i)}}{\partial X_{k}^{(i)}},\qquad\bm{\eta}_{k}^{(i)}=\frac{\partial J^{(i)}}{\partial\mathbf{Y}_{k}^{(i)}}.$ (14)
To connect this computational recurrence to the formal PMP framework, we must generalize it to an $\mathcal{F}_{t_{k}}$-adapted process $\lambda_{k}=\mathbb{E}[\lambda_{k}^{(i)}\mid\mathcal{F}_{t_{k}}]$. This transition from a path-specific value to an adapted process reveals a deep structural correspondence, which we formalize below.
Lemma 1 (Pathwise BPTT Recurrence).
Along any simulated trajectory $i$, the BPTT algorithm computes the backward recursion for the pathwise sensitivity $\lambda_{k}^{(i)}$ as:
@@
$\lambda_{k}^{(i)}$ $=\lambda_{k+1}^{(i)}+\lambda_{k+1}^{(i)}\bigl{[}r_{k}+\bm{\pi}_{k}^{\!\top}(\bm{\mu}_{k}-r_{k}\mathbf{1})\bigr{]}\Delta t+\bigl{[}e^{-\delta(t_{k}-t_{0})}U^{\prime}(C_{k}^{(i)})-\lambda_{k+1}^{(i)}\bigr{]}\frac{\partial C_{k}^{(i)}}{\partial X_{k}^{(i)}}\,\Delta t$
$\quad+\lambda_{k+1}^{(i)}\,\bm{\pi}_{k}^{\!\top}\bm{\sigma}_{k}\,\Delta\mathbf{W}_{k}^{X,(i)}.$ (15)
Proof.
The recurrence relation for the pathwise costate $\lambda_{k}^{(i)}:=\partial J^{(i)}/\partial X_{k}^{(i)}$ is derived by applying the chain rule. The total reward $J^{(i)}$ is decomposed into the utility from the current step $k$ and the total future reward $J_{k+1}^{(i)}$.
$\lambda_{k}^{(i)}=\frac{\partial}{\partial X_{k}^{(i)}}\left(e^{-\delta(t_{k}-t_{0})}U(C_{k}^{(i)})\Delta t\right)+\frac{\partial J_{k+1}^{(i)}}{\partial X_{k}^{(i)}}.$
The second term becomes $\lambda_{k+1}^{(i)}(\partial X_{k+1}^{(i)}/\partial X_{k}^{(i)})$. Differentiating the discrete wealth dynamics, which includes the consumption term $-C_{k}^{(i)}\Delta t$, with respect to $X_{k}^{(i)}$ yields:
$\frac{\partial X_{k+1}^{(i)}}{\partial X_{k}^{(i)}}=1+\bigl{(}r_{k}+\bm{\pi}_{k}^{\top}(\bm{\mu}_{k}-r_{k}\mathbf{1})\bigr{)}\Delta t-\frac{\partial C_{k}^{(i)}}{\partial X_{k}^{(i)}}\Delta t+\bm{\pi}_{k}^{\top}\bm{\sigma}_{k}\Delta\mathbf{W}_{k}^{X,(i)}.$
Substituting these components and rearranging the terms for $\lambda_{k}^{(i)}$ yields the expression in Eq. (1). ∎
Lemma 2 (Quadratic-Covariation Drift).
Let $\lambda_{k+1}^{(i)}=\lambda_{k+1}+\widetilde{\mathbf{Z}}^{\lambda X}_{k}\Delta\mathbf{W}_{k}^{X,(i)}+\widetilde{\mathbf{Z}}^{\lambda Y}_{k}\Delta\mathbf{W}_{k}^{Y,(i)}$ be the martingale representation of $\lambda_{k+1}^{(i)}$, where $\widetilde{\mathbf{Z}}^{\lambda X}_{k}$ is a row vector. The conditional expectation of the stochastic term in Eq. (1) is: